In the Recession Union, there are currently 2,500,000 people employed and 500,000 actively looking for work. In particular, of those unemployed, the government of the Recession Union calculates that one fourth of those who are unemployed owe their unemployment to the current economic downturn the country is experiencing. The rest are unemployed for different reasons. What is the natural rate of unemployment in the Recession Union? (Note for answering: If your answer is 20%, type into the answer box 20).
Answer #1

Those unemployed due to downturn are cyclically unemployed whose number os (500,000 / 4) = 125,000.

(Structurally + Frictionally) unemployed = 500,000 - 125,000 = 375,000

Natural rate of unemployment = (Structurally + Frictionally) unemployed / (Employed + Total unemployed)

= 375,000 / (2,500,000 + 500,000)

= 375,000 / 3,000,000

= 0.125

= 12.5 (%)